show me how to set the problem up. 175 cubic yds of pool water converting to weight in lbs of water
Answer by mananth(16946)

1cu. yard = 1yard * 1 yard * 1 yard
1 cu.yard = 3feet *3feet*3feet
1cu.yard = 9 cu.feet.
175 cu.yards = 175 * 9 = 1575 cu.ft.
1cu.ft of water = 62.4 pounds
1575 cu.ft = 1575*62.4
98,280 lbs.